the measures of 2 complementary angles are x and 2x + 12. find the measures of the angles.
Answer by nerdybill(7384)

the measures of 2 complementary angles are x and 2x + 12. find the measures of the angles.
.
If two angles are complementary, their sum totals 90 degrees:
.
x + 2x+12 = 90
3x+12 = 90
3x = 78
x = 26 degrees (first angle)
.
second angle:
2x+12 = 2(26)+12 = 52+12 = 64 degrees
.
answer: 26 and 64 degrees